A normally distributed population with 1000 observations has a mean of 100 and a standard devi-
ation 10. A sample of size 200 is randomly drawn from the population.
What is the standard error of the sampling distribution?
1
Expert's answer
2020-03-30T11:59:01-0400

The standard error of the mean (SEM) can be expressed as:


"SE={\sigma \over \sqrt{n}}={10 \over \sqrt{200}}={\sqrt{2} \over 2}\approx0.707"